# Information Loss under Coarse-Grained Partitions: A Discrete Framework for AI Ethics

### v6

@@ 10 $\frac{1}{10}$ $\frac{1}{11}$
However, this assumption is generally unrealistic: even among passing students, scores within the same category (grain) need not be evenly distributed. At the same time, once we coarse-grain the score scale, we want the coarsened distribution to preserve the total probability mass assigned to each grain, and we would like to define a canonical representative distribution that introduces no additional structure beyond this coarse information. To this end, we define Categorical Unification as follows.
Definition 5 (Categorical unification (CU) and $D_{\mathrm{KL\text{-}CU}}$).
Let $U$ be a finite set and let $\pi$ be a partition of $U$ with grain family $\mathfrak{G}_{\pi}=\{G_{\pi,i}\}_{i\in I_{\pi}}$. Let $P_{U}$ be a probability distribution on $U$.
(i) The coarse-grained distribution induced by $\pi$ is the pushforward $P_{\pi}$ on $I_{\pi}$ defined by
$P_{\pi}(i):=P_{U}(G_{\pi,i})=\sum_{u\in G_{\pi,i}}P_{U}(u),\qquad i\in I_{\pi}.$
Here, the category index $i$ represents the grain $G_{\pi,i}\subseteq U$; hence $P_{\pi}(i)$ is the total probability mass assigned to that grain, obtained by summing $P_{U}(u)$ over all $u\in G_{\pi,i}$.
(ii) The CU-unification of $P_{U}$ (through $\pi$) is the probability distribution $Q^{\mathrm{CU}}$ on $U$ defined by
Here, $Q^{\mathrm{CU}}$ spreads the coarse-grained mass $P_{\pi}(i)$ uniformly across the grain $G_{\pi,i}$; thus each $u\in G_{\pi,i}$ receives probability $Q^{\mathrm{CU}}(u)=P_{\pi}(i)/|G_{\pi,i}|$.
@@
where $D_{\mathrm{KL}}(P\|Q)=\sum_{u\in U}P(u)\log\!\frac{P(u)}{Q(u)}$ with the usual convention $0\log(0/q)=0$.
@@
The principle of maximum entropy provides a theoretical justification for CU. It states that among all possible probability distributions satisfying a given set of constraints, the one that maximizes entropy (i.e., uncertainty) is the least biased and most informative distribution [jaynes1957]. In the case of pass/fail classification, however, we are not only constrained by the requirement that probabilities sum to one but also by the fact that we know the number of students in each category. Specifically, there are six students in the fail category and four in the pass category, meaning that the total probability for each category is $\frac{6}{10}$ and $\frac{4}{10}$, respectively. According to the principle of maximum entropy, the probability distribution within each category should be uniform to maximize entropy. That is, within the fail category, the assumption that maximizes entropy (and therefore least biased) is a uniform distribution among the five elements (i.e., 0 to 4), and similarly, a uniform distribution should be assumed within the pass category (i.e., 5 to 10). Therefore, CU assigns uniform probability within each category, which corresponds to the entropy-maximizing distribution given the category-level constraints (Table 3). Since an entropy-maximizing distribution is the one that assumes the least additional information, CU can be considered the most unbiased method for probability assignment within coarse-grained categories.
@@
Let $U$ be a finite set and let $\pi$ be a partition of $U$ with grain family $\mathfrak{G}_{\pi}=\{G_{\pi,i}\}_{i\in I_{\pi}}$. Fix $P\in\Delta(U)$ and write
@@
Define the categorical unification (CU) of $P$ (relative to $\pi$) by
@@
$H(Q)=H(p)+\sum_{i\in I_{\pi}}p_{i}\,H(Q_{i}),$
where $H(p):=-\sum_{i}p_{i}\log p_{i}$ depends only on $p$ (hence is constant over $\mathcal{C}(p,\pi)$). Therefore, maximizing $H(Q)$ over $\mathcal{C}(p,\pi)$ reduces to maximizing each $H(Q_{i})$ over the simplex $\Delta(G_{\pi,i})$. For a fixed finite set $G_{\pi,i}$, Shannon entropy is maximized uniquely by the uniform distribution on $G_{\pi,i}$, i.e., $Q_{i}(x)=1/|G_{\pi,i}|$ for all $x\in G_{\pi,i}$. Thus, for all $x\in G_{\pi,i}$,
@@ 10 $\frac{1}{10}$ Pass $\frac{1}{15}$
Let $U$ be a finite set and let $\pi$ be a partition of $U$ with grain family $\mathfrak{G}_{\pi}=\{G_{\pi,i}\}_{i\in I_{\pi}}$. Let $P_{U}$ be a probability distribution on $U$, and define the induced coarse-grained distribution
@@
Equivalently, for every $i\in I_{\pi}$ with $P_{\pi}(i)>0$, the conditional distribution of $P_{U}$ given the event $\{u\in G_{\pi,i}\}$ is uniform on $G_{\pi,i}$:
$P_{U}(\,\cdot\,\mid G_{\pi,i})=\mathrm{Unif}(G_{\pi,i}).$
@@
4.4 Minimizing Information Loss
We aim to make the probability distribution of the coarse-grained pass/fail evaluation as similar as possible to the original probability distribution of scores ranging from 0 to 10. Here, “similar” specifically refers to minimizing information loss, which is quantitatively measured by minimizing the $D_{\text{KL-CU}}$. Suppose we need to divide those ten students (Table 1) into pass and fail groups, but we can determine the passing threshold $(T)$. What score should be used as $T$?
A seemingly natural approach is to use the average score. Since the mean score in this test is 4.1, setting the passing threshold at $T=4$ may seem reasonable. If we classify students scoring below 4 as failing and those scoring 4 or above as passing, we obtain a $D_{\text{KL-CU}}$ of approximately 0.524, which is relatively low. However, this is not the minimum. The lowest $D_{\text{KL-CU}}$ is achieved when the passing threshold is set at $T=7$, resulting in a value of 0.381 (Table 4, Figure 2).
@@ 4.4 Minimizing Information Loss
If an AI teacher performed this calculation and told a human teacher, “Based on this test, only $s_{5}$ should pass”, the human teacher might find it counterintuitive. However, its interpretation is context-dependent. If the number of passing students is fixed (e.g., a predetermined quota of three students), minimizing $D_{\text{KL-CU}}$ may not always be the correct objective. In such cases, the goal is not to preserve the original probability distribution but rather to adhere to external constraints, such as ensuring that exactly three students pass, regardless of their relative performance. This suggests that setting a fixed pass quota can significantly distort the original probability distribution.
Many educators may have encountered situations in which students had widely varying abilities or had to pass students whose performance was lower than they would ideally consider acceptable. This raises a critical question: If the information loss is excessively large, should AI provide an advisory to the teacher? To address this, it may be beneficial to equip the AI teacher with a feature that advises, “The school has mandated that three students must pass this test. Do you believe this number is truly appropriate?” This would allow for human oversight in cases where strict adherence to an imposed quota might compromise fairness or accuracy.
In some contexts, setting the threshold at $T=7$, where only $s_{5}$ passes, may actually align with intuition. Consider a scenario in which the goal is to divide the class into two distinct groups: one for mathematically advanced students (e.g., an honors class) and another for those who need more support. The original score distribution implies that only one student significantly outperforms the others in mathematics; accordingly, $T=7$ identifies a single clear outlier ($s_{5}$). Thus, setting $T=7$ produces a classification that best preserves the structure of the original probability distribution while maintaining a meaningful distinction between high and low performers.
This analysis highlights how the calculation of $D_{\text{KL-CU}}$ can serve as a bridge between AI-driven decision-making and human intuition, offering a foundational example for both XAI and CE.
@@ 5.1 Mathematical Foundation of Coarse Ethics
CGPs have demonstrated that the rationality of coarse evaluations, as argued by CE, can be rigorously analyzed. The coarse evaluations we use in daily life—for instance, the assessment that “Olivia received an excellent grade on her mathematics test”—exhibit a structured mathematical form within CGPs. The fact that an evaluation is coarse—for example, that Olivia’s specific score difference from other students who also received an “excellent” rating is intentionally disregarded—does not imply that the evaluation is ambiguous. Here, we can clearly distinguish between coarseness and ambiguity in assessment.
Traditional ethical theories, particularly consequentialist (e.g., felicific calculus by Bentham [bentham1789introduction]) and virtue-based approaches (e.g., Stoicism), often emphasize continuity in moral evaluation, where actions that align more closely with a given ideal are ranked proportionally higher. However, CE suggests that even when individuals differ significantly in their alignment with an ideal, they may still be evaluated equivalently [izumoweng2022]. This aligns with Immanuel Kant’s criticism of excessive ethical precision, particularly his rejection of casuistry—the attempt to make moral laws overly specific through case-by-case analysis [kant2017]. A simple example illustrates this point: Suppose a student receives a failing grade of 30 points on an essay from a human teacher. Dissatisfied, they submit the same essay to an AI teacher for evaluation, which assigns them a score of 30.1 while their peer, who was also given 30 by the human teacher, receives 29.8 from the AI. If the student then argues that the AI’s refined scoring justifies a reconsideration of their grade, such a claim would be unreasonable. From the teacher’s perspective, both evaluations reflect insufficient performance, and fixating on a marginal difference of 0.3 points does not meaningfully change the outcome.
5.2 Tradeoff Between Evaluation Efficiency and Information Retention
As demonstrated in this study, when $Q(i)$ in $D_{\text{KL}}$ is processed through categorical unification, information is preserved only in specific cases. Our findings suggest a potential trade-off between evaluation efficiency and information retention. While this study does not provide a formal proof, we hypothesize that as evaluation processes prioritize simplicity and accessibility, the loss of fine-grained details becomes inevitable. Further research could investigate whether there exists a Pareto-optimal balance between evaluation efficiency and retained information in AI-driven decision-making.
Assuming a uniform distribution within a category is the most reasonable default when no additional information is available. When only a pass label is recorded, any distinction between different passing scores is inherently erased, making a categorically uniform distribution the simplest assumption that reflects this lack of distinguishing data. While alternative assumptions, such as normal or skewed distributions, might seem more refined, they require external data that is often unavailable in real-world grading systems. If historical trends suggest a particular distribution (e.g., most passing students score near the middle of the range), Bayesian updates could provide a better-informed probability model. However, in the absence of such priors, assuming a categorically uniform distribution remains the least biased approach.
5.3 Application to Explainable AI
We apply the above discussion to the challenges identified in Section 1.2. One key application is in refining AI-driven safety advisories. For example, in autonomous driving systems, the coarse-graining of safety warnings must not result in the loss of critical risk information. Instead of issuing a generic warning such as “Your driving is dangerous”, CGPs can guide AI systems to adaptively determine the level of warning granularity, ensuring that important details about immediate risks (e.g., speed excess relative to road conditions or proximity to nearby obstacles) are retained within the coarsened message. This approach allows for a structured balance between reducing cognitive overload and preserving necessary safety information.
Additionally, CGPs offer a refined approach to addressing the limitations of the three-tiered explanatory model in CE. Rather than strictly categorizing explanations into those for experts, general users, and individuals with cognitive difficulties, CGPs allow for dynamic adaptation of explanation granularity by integrating user profiling mechanisms. In a medical diagnostic AI, for example, an expert might receive a detailed probability breakdown of differential diagnoses, while a general patient would be presented with a high-level summary emphasizing key risks and actionable next steps. This ensures that explanations remain accessible without sacrificing necessary details for informed decision-making. This personalized granularity adjustment can improve the effectiveness of AI explanations in domains such as medical diagnostics, legal decision-making, and financial advising, where different users require distinct levels of detail in their assessments.

### v7

@@ 10 $\frac{1}{10}$ $\frac{1}{11}$
However, this assumption is generally unrealistic: even among passing students, scores within the same category (grain) need not be evenly distributed. At the same time, once we coarse-grain the score scale, we want the coarsened distribution to preserve the total probability mass assigned to each grain, and we would like to define a canonical representative distribution that introduces no additional structure beyond this coarse information. To this end, we define Categorical Unification (CU) as follows.
Definition 5 (Categorical unification and $D_{\mathrm{KL\text{-}CU}}$).
Given the underlying scale $(U,\leq)$, let $P_{U}$ be a probability distribution on $(U,2^{U})$.
(i) The coarse-grained distribution induced by $\pi$ is the probability distribution $P_{\pi}$ on $I_{\pi}$ defined by
$P_{\pi}(\{i\}):=P_{U}(G_{\pi,i})=\sum_{u\in G_{\pi,i}}P_{U}(\{u\}),\qquad i\in I_{\pi}.$
Equivalently, in the discrete setting we write $P_{U}(u):=P_{U}(\{u\})$ and $P_{\pi}(i):=P_{\pi}(\{i\})$.
(ii) The CU-unification of $P_{U}$ (via $\pi$) is the probability distribution $Q^{\mathrm{CU}}$ on $U$ defined by
$Q^{\mathrm{CU}}(u):=\frac{P_{\pi}(\{i\})}{|G_{\pi,i}|}\qquad\text{for }u\in G_{\pi,i}.$
Since $\pi$ is a partition, each $u\in U$ belongs to a unique grain $G_{\pi,i}$, so $Q^{\mathrm{CU}}$ is well-defined. Note that $Q^{\mathrm{CU}}$ is indeed a probability distribution on $U$: it redistributes each coarse mass $P_{\pi}(\{i\})$ uniformly over the points $u\in G_{\pi,i}$. In particular,
$\sum_{u\in U}Q^{\mathrm{CU}}(u)=\sum_{i\in I_{\pi}}\sum_{u\in G_{\pi,i}}\frac{P_{\pi}(\{i\})}{|G_{\pi,i}|}=\sum_{i\in I_{\pi}}P_{\pi}(\{i\})=1.$
@@
where
$D_{\mathrm{KL}}(P\|Q)=\sum_{u\in U}P(u)\log\!\frac{P(u)}{Q(u)},$
with the convention that any term with $P(u)=0$ is taken to be 0.
The point of Definition 5 is not to compare the original distribution directly with the coarse representation itself. Rather, $D_{\mathrm{KL\text{-}CU}}$ measures how far the original distribution deviates from a canonical fine-scale reconstruction induced by the coarse representation under minimal assumptions.
@@
Additionally, the principle of maximum entropy provides a theoretical justification for CU. It states that among all probability distributions satisfying a given set of constraints, the one that maximizes entropy (i.e., uncertainty) is the least biased and least informative beyond the imposed constraints [jaynes1957]. In the case of pass/fail classification, however, we are not only constrained by the requirement that probabilities sum to one but also by the fact that we know the number of students in each category. Specifically, there are six students in the fail category and four in the pass category, meaning that the total probability for each category is $\frac{6}{10}$ and $\frac{4}{10}$, respectively. According to the principle of maximum entropy, the probability distribution within each category should be uniform to maximize entropy. That is, within the fail category, the assumption that maximizes entropy (and therefore least biased) is a uniform distribution among the five elements (i.e., 0 to 4), and similarly, a uniform distribution should be assumed within the pass category (i.e., 5 to 10). Therefore, CU assigns uniform probability within each category, which corresponds to the entropy-maximizing distribution given the category-level constraints (Table 3). Since an entropy-maximizing distribution is the one that assumes the least additional information, CU can be considered the most unbiased method for probability assignment within coarse-grained categories.
@@
Given the underlying scale $(U,\leq)$, fix $P\in\Delta(U)$ and write
@@
Define the CU of $P$ (relative to $\pi$) by
@@
The constraint set $\mathcal{C}(p,\pi)$ fixes only the coarse information, namely the total mass $p_{i}$ assigned to each grain $G_{\pi,i}$, while leaving the distribution within each grain free. For example, if a grain is $G_{\pi,i}=\{8,9,10\}$ and $p_{i}=0.3$, then any assignment of nonnegative probabilities $(Q(8),Q(9),Q(10))$ with $Q(8)+Q(9)+Q(10)=0.3$ is admissible, such as $(0.1,0.1,0.1)$ or $(0.29,0.01,0)$. Proposition 2 shows that, among all fine-grained distributions consistent with these coarse masses, the CU distribution $P^{\mathrm{CU}}$ is the least informative one in the sense of maximum Shannon entropy (in the above example, this corresponds to the uniform assignment $(0.1,0.1,0.1)$ on $G_{\pi,i}=\{8,9,10\}$): it preserves the grain-level totals $p_{i}$ and otherwise spreads mass uniformly within each grain. In particular, CU can be viewed as the canonical “uninformative” refinement of the coarse-grained distribution back to the original scale $U$.
@@
$H(Q)=H(p)+\sum_{i\in I_{\pi}}p_{i}\,H(Q_{i}).$
This identity shows that the total entropy of $Q$ consists of the entropy $H(p)$ at the coarse-grained level (which grain is selected) plus the average within-grain entropy $\sum_{i\in I_{\pi}}p_{i}H(Q_{i})$, weighted by the grain masses $p_{i}$. $H(p):=-\sum_{i}p_{i}\log p_{i}$ depends only on $p$ (hence is constant over $\mathcal{C}(p,\pi)$) Therefore, maximizing $H(Q)$ over $\mathcal{C}(p,\pi)$ reduces to maximizing each $H(Q_{i})$ over the simplex $\Delta(G_{\pi,i})$. For a fixed finite set $G_{\pi,i}$, Shannon entropy is maximized uniquely by the uniform distribution on $G_{\pi,i}$, i.e., $Q_{i}(x)=1/|G_{\pi,i}|$ for all $x\in G_{\pi,i}$. Thus, for all $x\in G_{\pi,i}$,
@@ 10 $\frac{1}{10}$ Pass $\frac{1}{15}$
Given the underlying scale $(U,\leq)$, let $P_{U}$ be a probability distribution on $U$, and define the induced coarse-grained distribution
@@
Equivalently, for every $i\in I_{\pi}$ with $P_{\pi}(i)>0$ and every $A\subseteq G_{\pi,i}$,
$P_{U}(A\mid G_{\pi,i})=\frac{|A|}{|G_{\pi,i}|}.$
@@
This theorem characterizes the situation in which elements that are mapped to the same grain (that is, the same category) after coarsening are already assigned uniform probability mass within that grain under the pre-coarsening distribution $P_{U}$.
4.3 Optimizing Information Loss
4.3.1 Minimizing the KL Divergence
When moving from simple examples such as pass/fail classification to the general design of coarse-grained evaluations, an additional question arises: how should we choose among multiple possible coarse-grainings? Once admissible coarse-grainings have been restricted to CGPs satisfying the structural requirements of CE, the remaining question is how to compare multiple admissible partitions. The role of $D_{\mathrm{{KL\text{-}CU}}}$ in this section is to provide one such comparative criterion.
As the zero-loss theorem for $D_{\mathrm{KL\text{-}CU}}$ shows, making information loss vanish in the sense of this measure requires that the original fine-grained distribution already be uniform within each grain. In practice, however, the criteria for coarsening an evaluation are rarely chosen so as to align with the original distribution in this way. For example, in grading a mathematics test, a teacher would not normally create a category simply because the numbers of students scoring 61, 62, and 63 happen to be equal. The zero-loss theorem therefore should not be understood as providing a practical guideline for optimizing categorization. Rather, it shows that practical categorization is usually guided by considerations other than the complete elimination of information loss in the sense of $D_{\mathrm{KL\text{-}CU}}$.
One example in which zero information loss cannot be expected is when the number of categories is fixed in advance—for instance, when outcomes must be classified only as pass or fail. One possible idea, then, is to give up on achieving zero information loss and instead set the threshold so as to minimize information loss. Suppose we need to divide those ten students (Table 1) into pass and fail groups, but we can determine the passing threshold $(T)$. What score should be used as $T$? In this setting, when the number of grains is fixed, minimizing information loss provides a natural yet straightforward way to formulate the objective function. The lowest $D_{\text{KL-CU}}$ is achieved when the passing threshold is set at $T=7$, resulting in a value of 0.381 (Table 4, Figure 2).
@@ 4.3.1 Minimizing the KL Divergence
However, reducing the coarse-graining design problem to mere information-loss minimization is not always adequate. In many settings (e.g., pass/fail decisions), the relevant objective is not to reproduce the original score distribution, but rather to minimize decision costs (such as false positives and false negatives) or to optimize expected utility. Consider again the arithmetic test example above. If the pass threshold $T$ is chosen so as to minimize $D_{\mathrm{KL\text{-}CU}}$, the optimum may be a threshold such as $T=7$, under which only the student $s_{5}$ passes. This is reasonable from the viewpoint of distributional fidelity, because it preserves, within the coarse pass/fail representation, the fact that $s_{5}$ stands out as an exceptionally strong performer in this class of ten students.
Yet such an information-loss-minimization principle is justified only when the teacher’s primary aim is to preserve information about the original score distribution. In educational practice, pass/fail evaluations are often tied instead to an operational criterion, such as whether a student can keep up with the next course module. For example, if scoring at least 6 points is sufficient to follow the next class, the teacher may reasonably wish to set the pass threshold at $T=6$. In that case, selecting $T=7$ by minimizing $D_{\mathrm{KL\text{-}CU}}$ may still be meaningful from the viewpoint of distributional fidelity, but it fails to reflect the actual decision objective and therefore becomes an inappropriate design principle for the coarse-graining.
4.3.2 An Optimization View of Coarse-Grained Design
Therefore, KL-based information-loss minimization should be understood as a principled baseline objective, one that may need to be supplemented or replaced when decision-theoretic or normative constraints are dominant. Formally, given a finite totally ordered set $(U,\leq)$ and a distribution $P$ on $U$, we consider the optimization problem
$\min_{\pi\in\mathcal{P}_{\mathrm{int}}(U,\leq)}\;\;D_{\mathrm{KL}}\!\left(P\,\middle\|\,Q^{\mathrm{CU}}_{\pi}\right)\;+\;\lambda\,\Omega(\pi),$ (2)
where $\mathcal{P}_{\mathrm{int}}(U,\leq)$ denotes the collection of interval coarse-grained partitions, $Q^{\mathrm{CU}}_{\pi}$ is the CU-unification associated with $\pi$, $\Omega(\pi)$ is a complexity penalty (e.g., the number of grains $m(\pi)$ or a description-length proxy), and $\lambda\geq 0$ controls the trade-off between informational fidelity and simplicity. In this formulation, coarse-graining is not merely an evaluative act, but also the choice of a representation guided by optimization, and one that can be adapted to domain-specific constraints and interpretability demands.
The case $\lambda=0$ corresponds to the limiting situation in which minimizing $D_{\mathrm{KL\text{-}CU}}$ is itself the sole objective. In that case, the finest partition is trivially optimal, since no penalty is imposed on excessive refinement. The present framework is intended primarily for the nontrivial case $\lambda>0$, where the design of coarse-graining requires a trade-off between informational fidelity and coarsening cost. From this perspective, the role of $\lambda$ is not merely technical, but expresses the practical fact that coarse-grained evaluation is introduced precisely because interpretability, simplicity, and other costs of fine-grained representation cannot be ignored.
We leave a systematic study of (2), including computational aspects and constraint variants, to future work.
@@ 5 Discussion
We show that the core ideas of CE can be given a precise mathematical formulation. Coarse evaluations in everyday life—for example, the statement that “Olivia received an excellent grade on her mathematics test”—exhibit a clear mathematical structure within the framework of coarse-grained partitions. The fact that an evaluation is coarse—for instance, that Olivia’s exact score difference from other students who also received an “excellent” rating is intentionally ignored—does not mean that the evaluation is ambiguous. In this sense, CGPs make it possible to distinguish clearly between coarseness and ambiguity in assessment. Therefore, CE’s attempt to distinguish coarse evaluations from arbitrary or haphazard evaluations can be justified mathematically.
CE has proposed, as a criterion for comparing fine-grained and coarse-grained evaluations, that a coarse-grained evaluation should not reverse the value ordering embodied in the fine-grained one. However, as the previous literature itself acknowledges, multiple coarse-grained evaluations can satisfy this condition for the same fine-grained evaluation. The problem, therefore, is that this criterion alone cannot sufficiently determine how the coarsening should be carried out. In response, we have proposed a new criterion based on comparing the probability distributions associated with the fine-grained and coarse-grained evaluations, and on minimizing the sum of information loss and coarsening cost. When the fine-grained evaluation is given as a finite discrete totally ordered set, a coarse-grained evaluation constructed within the framework of CGPs allows the corresponding probability distributions and information loss to be defined and computed rigorously. Moreover, the central result of this paper—the zero-loss theorem for $D_{\mathrm{KL\text{-}CU}}$—shows that coarsening yields zero information loss in the sense of this measure only when the original fine-grained distribution is already uniform within each grain. Hence, coarse-grained evaluations with zero information loss, in the sense of $D_{\mathrm{KL\text{-}CU}}$, occur only in highly exceptional cases in actual evaluative practice.
These observations are also relevant to explainable AI. In many AI systems, the model first produces a fine-grained internal evaluation of its target. This is true, for example, in AI-assisted driving systems that compute detailed risk scores, in medical AI that estimates diagnostic likelihoods, and in hiring or interview AI that generates fine-grained personnel assessments. From this perspective, one important problem in XAI is how such fine-grained internal evaluations should be transformed into coarser outputs that are intelligible and usable for human agents. A simple example is AI-assisted driving. Suppose that the system internally represents driving risk on a fine-grained scale such as $U=\{0,1,\dots,100\}$, while the human-facing interface must display only a small number of warnings, such as safe, caution, and danger. In this case, a CGP represents the transformation from the fine-grained risk scale to these coarse warning categories. The corresponding quantity $D_{\mathrm{KL\text{-}CU}}$ then measures how far the original fine-grained risk distribution deviates from the canonical fine-scale reconstruction induced by that coarse warning system under minimal assumptions. This makes the design problem more precise. The issue is not only whether the warning system remains intelligible to the driver, but also how much fine-grained risk information is lost when the internal evaluation is compressed into a small number of human-interpretable categories. In such cases, the framework proposed in this paper becomes directly relevant: a principled warning system should balance informational fidelity against the coarsening cost associated with human cognitive burden, instruction complexity, and response usability.